OpenStudy (anonymous):

let f(x)= –5x – 4 and g(x) = 6x – 7. Find f(x) + g(x). –11x + 3 x + 3 –11x – 11 x – 11 I think its D. , but not sure ?

OpenStudy (anonymous):

just add them –5x – 4 + 6x – 7 x - 11 is correct

OpenStudy (anonymous):

Thanks , that's what I thought !! (:

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

Can you help me w/ a different one please ?

OpenStudy (anonymous):

5 bucks :P, yes

OpenStudy (anonymous):

let f(x) = –5x + 3 and g(x) = 6x – 2. Find f • g and its domain. –30x^2 + 28x – 6; all real numbers except x = 1/3 –15x^2 – 8x – 12; all real numbers except x = 3/5 –30x^2 + 28x – 6; all real numbers –15x^2 – 8x – 12; all real numbers

OpenStudy (anonymous):

I think its C. , but not sure && ok 5 $ on your way (;

OpenStudy (anonymous):

f • g =(–5x + 3)(6x – 2) –30x^2 + 28x – 6; all real numbers

OpenStudy (anonymous):

checked the mail there is nothing :p

OpenStudy (anonymous):

I was righttt (: thanksss & its gonna take a few days silly . lol (:

OpenStudy (anonymous):

anytime